# Fairshare Data Pricing via Data Valuation for Large Language Models

### v3

@@ 1 Introduction
High-quality training data is foundational to building effective and reliable large language models (LLMs). As LLMs take on increasingly complex tasks today – such as coding[1], reasoning [2], and AI4Science [3] – they rely heavily on carefully curated, human-annotated data. This growing demand has triggered a "generative data gold rush", with major tech companies racing to acquire training data, fueling the rise of a nascent AI data market [4]. In this market, AI firms create networks of short-term contract workers to generate data labels, resembling an Uber-like gig economy for data [4].
However, the current AI data market operates with limited oversight and is widely criticized for a lack of transparency and fairness in pricing [5, 6, 7, 8]. Data prices are largely low and fail to reflect the quality or effort involved, threatening the sustainability and quality of data supply [9, 10, 11]. In particular, for data sellers, such as human annotators or content creators, the prevailing market routinely undervalues their labor, offering compensation that neglects the skill, effort, and downstream value of their contributions. These harms are especially concentrated in low-wage labor markets, where annotators often face overwork, underpayment, and exclusion from decision-making [12]. This reflects a broader ethical concern known as “AI parachuting”, where developers extract data from marginalized communities [13, 14, 15], tieing to wider debates on epistemic injustice and data colonialism [16, 17, 18]. As one civil rights advocate noted on 60 Minutes, “They don’t pay well… they could pay whatever, and have whatever working conditions.” [^1] CBS, 60 Minutes
Motivated by these issues, we present a fair pricing framework for the LLM training data market to promote equitable and sustainable generative AI ecosystems. Economic theory suggests that prices should reflect the value delivered to the buyer – signaling quality and alignment [19]. Guided by this, our framework introduces fairshare pricing based on established data valuation techniques for LLMs [20, 21, 22, 23], which quantify each dataset’s contribution to model performance. Buyers and sellers both have access to data valuation scores, which guide decision-making on both sides: buyers use them to select datasets under budget constraints to maximize utility (i.e., a standard measure of welfare or satisfaction [24, 25]), while sellers use them to set prices based on the anticipated demand from buyers. This shared access – enabled by our assumption of information transparency – supports procedural fairness [26], improves seller participation, and enhances overall data quality.
@@
Second, we empirically validate our approach through simulations of buyer-seller interactions in data markets. We focus on training open-source LLMs on complex NLP tasks, including math problems [27], medical diagnosis [28], and physical reasoning [29]. Analyzing both pricing and valuation outcomes, we find that under fairshare pricing, buyers achieve higher model performance per dollar spent, making it particularly beneficial for those with limited budgets. In addition, our simulations of long-term market dynamics demonstrate that fairshare pricing encourages sustained seller participation, resulting in a stable and sufficient supply of training data over time compared to exploitative pricing. These findings show that our framework’s data-valuation-based pricing not only improves short-term training efficiency, but also ensures the long-term viability of the data market.
Finally, to evaluate the robustness and method-agnostic applicability of our framework, we conduct an ablation study using a diverse set of data valuation methods (including BM25 [30], $\text{Infl}_{\text{IP}}$ [22], and Datainf[21]) – selected for their scalability and efficiency in LLM tasks. Across all variants, our fairshare pricing framework consistently delivers mutually beneficial results for both buyers and sellers in the LLM data market, confirming that its performance is not tied to any specific data valuation technique. This analysis underscores a key strength of our approach: buyers and sellers can flexibly choose valuation methods tailored to their downstream needs without compromising the incentive-aligned structure of the market, and demonstrates the broad applicability of our solution.
@@ 2 Related Work
Data Market Design and Exploitative Pricing. Recent studies on ML data markets focus on platform-based pricing and allocation mechanisms such as auctions and personalized pricing [31, 32, 33, 34, 35]. These frameworks typically aim to maximize profit or efficiency, assuming myopic buyers and fixed seller participation. They also consider information transparency ensures that sellers have access to the same data valuation scores used by buyers [36, 33]. The study was mainly conducted on classic machine learning models rather than large language models.
@@
Concurrently, research has highlighted ethical concerns in LLM data acquisition, particularly around exploitative pricing. Annotation work – central to LLM training – is often outsourced to low-wage workers with minimal labor protections [9, 10, 37, 38], who are frequently exposed to harmful content [5, 12, 39]. Several studies document low pay, precarious labor, and lack of protections in global data supply chains [16, 17, 18]. Although these studies highlight systemic exploitation in data labor, their findings have yet to be integrated into formal models of LLM data pricing. Economic theory suggests that misaligned compensation undermines efficiency, discourages participation, and lowers data quality. In LLM contexts, this reduces data availability and degrades model performance [40, 41, 42], underscoring the need for value-aligned pricing mechanisms.
@@
Game-Theoretic Models of Pricing and Participation. In many domains, pricing problems are modeled thorugh game theory. Particularly, those involving decision-making are frequently modeled using Stackelberg games and bilevel optimization [43]. Although these frameworks have not yet been applied to LLM data market, they have been extensively used in contexts such as resource allocations [44], energy market [45, 46], supply chains [47], distributed systems [48]. In these settings, agents typically optimize revenue, efficiency, or utility while anticipating the strategic responses of others. In dynamic settings, these models extend to repeated interactions, where players balance immediate gains against long-term benefits [49, 43]. In repeated interactions, fairness perceptions are critical: studies show that perceived pricing unfairness can erode trust and reduce participation [50, 51, 52, 53].
@@
Data Valuation Methods: Data valuation estimates the contribution of training examples to model performance [54]. Examples include influence-function-based methods [55], which estimate data utility via model gradient computations. Variations such of this method enhance efficiency by approximating/bypassing the inverse-Hessian [56], or utilizing lower dimension model gradients (e.g., Datainf [21]), which balance efficiency and accuracy, and make them suitable for the LLM realm. Beyond the influence of loss functions, previous research [57] has examined data valuation with respect to fractional utility functions.
@@
Data valuation approaches are effective for LLM training data selection [22, 23, 58] and other applications including toxicity detection [59], memorization analysis [60], training optimization [61], and and active sampling [62]. Shapley-based method is another direction [63, 64, 65, 36], such as CS-Shaply[66], In-run Shapley[67], DU-Shapley[68], etc. Simpler methods like BM25 [30] offer a model-agnostic baseline based on lexical similarity[69, 70]. While exact data value estimation is computationally costly at LLM scale, recent studies show that influence-based approximations correlate meaningfully with actual performance outcomes [71, 72]. For example, Jiao et al. [71] shows that influence-based methods have high correlation between their estimated data valuation scores and the oracle value, which is obtained through model re-training.
@@ 3.1 Data Market Definition
We formalize the LLM data market by modeling the decision-making processes of sellers and buyers in a non-cooperative Stackelberg game [35, 45, 48]. In this setup, sellers first set prices by anticipating buyer demand. Buyers then respond by selecting datasets to maximize their utility. [^2] Utility is defined as a general economic concept that measures an individual’s welfare, benefit, or satisfaction [24, 25], for example, monetary value, model fairness, or other domain-specific benefits. Each player makes decisions based solely on the current state of the market, without considering future updates or interactions.
Notably, our choice of a non-cooperative Stackelberg game fits the LLM data market’s reality: many independent, self-interested transactions where binding grand coalitions are infeasible, such as gig-work platforms [18]. In contrast, cooperative bargaining targets few-party, contractible surplus splits, which is a different regime and question. In this non-cooperative game, we model fairness behaviorally via seller participation/exit that keeps the model tractable, and the leader–follower timing mirrors real price setting and demand response.
@@
Following prior work [73], we assume information transparency: all participants observe each dataset’s value. Practically, this can be estimated using data valuation methods that quantify its contribution to LLM performance. These valuation scores – potentially generated by the platform or a trusted third party – serve as common signals informing both pricing and purchasing decisions.
@@
The value $v_{k}$ can be estimated using a data valuation method (e.g., Influence Function [55], DataInf [21], etc.). Once the data valuation score $v_{k}$ is estimated, the corresponding utility gain $u_{k}$ can be derived as a function of $v_{k}$. We use a task-specific mapping between data value $v_{k}$ and utility $u_{k}:v_{k}\rightarrow\mathbb{R}_{\geq 0}$. Appendix˜C lists common mappings between $u_{k}$ and $v_{k}$ in downstream tasks.
@@ 3.2 Exploitative Pricing: A Lose-Lose Outcome
Assumption˜1 models how the seller respond to exploitative pricing over time. It states that when offered a price below its ideal level, the seller become less likely to stay in the market, consistent with prior studies [74, 75]. For any exploitative pricing, the participation probability $P_{T}$ declines multiplicatively, leading to a sustained and irreversible drop in engagement.
@@
Assumption˜2 assumes the buyer values both immediate and future utility gains from acquiring training data. Major LLM developers invest in large-scale data acquisition and model training in expectation of long-term gains in performance, deployment value, and commercial returns [76, 77, 78, 79].
@@
Buyer’s Objective. With the previous setup, the buyer aims to maximize expected cumulative utility over an infinite horizon. Let $G(u_{t},b_{t})$ denote this value function, representing the buyer’s expected total cumulative utility at time $t$, conditioned on current utility $u_{t}$ and budget $b_{t}$. Then this objective satisfies the following Bellman equation[80]:
@@
In addition to our theoretical findings, we run a simplified dynamic market simulation with one buyer and one seller, making sequential decisions. The utility of the dataset and buyer’s budget varies randomly over time, and seller participation follows $\pi(p_{t},p_{t}^{*})=p_{t}/p_{t}^{*}$. Figures˜2(a) and 2(b) provides consistent evidence that exploitative pricing causes rapid seller exit and an immediate shortage of training data. This dynamic closely resembles the classic “market for lemons” problem, where underpricing drives out high-quality supply, ultimately leading to market collapse [8].
@@ 3.3 Fairshare Pricing: A Win-Win for Sellers and Buyers
And $t^{*}$ is increasing as $\delta$ increases.
@@ 4.1 Experimental Setup
Buyers and Models. We consider three buyers, each using a standard open-source LLMs: Llama-3.2-Instruct-1b [81], Pythia-1b, and Pythia-410m [82]. These models are pre-trained on different corpora and exhibit varying preferences for downstream post-training data [83].
Sellers and Datasets. We focus on challenging, human-annotated tasks: MathQA and GSM8K [27, 84] for math, MedQA [28] for medical diagnosis, and PIQA [29] for physical reasoning [85, 86, 87]. Table˜1 in Appendix˜F shows dataset splits and examples. We use the training splits as seller data and simulate the market dynamics from Section˜3, treating each task as a market scenario.
@@ 4.2 LLM Data Market Experiments
Exploitative Pricing Leads to Lose-Lose Outcomes: The exploitative pricing method, which sets uniformly low, fixed prices to reflect real-world practices ([11]), offers LLM developers short-term utility gains (see Figures˜4(a) and 4(b)). However, this approach systematically undervalues datasets and unfairly compensates annotators. As a result, sellers exit the market over time, triggering a collapse in data supply. Even well-funded LLM developers are unable to source sufficient data – hindering the advancement of LLMs.
@@ 4.3 Ablation Study: Effect of Different Data Valuation Methods
Data Valuation Methods. We consider the following methods, where each assigns a value to every data sample: (1) Constant baseline – assigns the same value, mimicking flat-rate pricing on platforms; (2) Random baseline – values drawn uniformly from $[0,1]$; (3) Semantic – uses BM25 [30] to compute average similarity to the representative set; (4) Influence-based: returns a score which leverages learning gradients to estimate a data sample’s avg. contribution to model learning of a representative dataset. Specifically, we use $\text{Infl}_{\text{IP}}$ [56, 22] and DataInf  [21], which are influence-based methods adapted for the LLM realm. See Appendix˜A for additional details.
Furthermore, we adopt a one-step training approach for data valuation, where the value of each data sample is estimated by performing a single training step on it and measuring the resulting change in model performance relative to the original model. Results are presented in in Section˜F.2. This approach serves as an “oracle” for influence-based methods [23, 56], as it directly quantifies data value through model training, although it is more computationally expensive than the previous listed methods.
@@
We also run the error analysis of data valuation methods. Following previous studies [71, 72], we compute the correlation between Oracle and $\text{Infl}_{\text{IP}}$, reporting spearman correlation of $0.54,0.42$ for MathQA and PIQA (see Figure˜8 in Appendix˜F). This underscores the opportunity for future advances in data valuation accuracy and scalability, which can be seamlessly integrated into our flexible fairshare framework.

### v4

@@ 1 Introduction
High-quality training data is foundational to building effective and reliable large language models (LLMs). As LLMs take on increasingly complex tasks today – such as coding[chen2021humaneval], reasoning [wei2022chain], and AI4Science [taylor2022galactica] – they rely heavily on carefully curated, human-annotated data. This growing demand has triggered a "generative data gold rush", with major tech companies racing to acquire training data, fueling the rise of a nascent AI data market [reuters2024bigtech]. In this market, AI firms create networks of short-term contract workers to generate data labels, resembling an Uber-like gig economy for data [reuters2024bigtech].
However, the current AI data market operates with limited oversight and is widely criticized for a lack of transparency and fairness in pricing [paul2024bigtech, zhang2024surveydatamarkets, metcalf2021algorithmic, akerlof1978market]. Data prices are largely low and fail to reflect the quality or effort involved, threatening the sustainability and quality of data supply [mason2009financialcrowds, hara2018datadrivenamt, cbsnews2024]. In particular, for data sellers, such as human annotators or content creators, the prevailing market routinely undervalues their labor, offering compensation that neglects the skill, effort, and downstream value of their contributions. These harms are especially concentrated in low-wage labor markets, where annotators often face overwork, underpayment, and exclusion from decision-making [time2023kenya]. This reflects a broader ethical concern known as “AI parachuting”, where developers extract data from marginalized communities [jo2020archives, birhane2021impossibility, qz_africa_parachute], tieing to wider debates on epistemic injustice and data colonialism [paullada2021datadiscontents, sambasivan2021datawork, gray2019ghost]. As one civil rights advocate noted on 60 Minutes, “They don’t pay well… they could pay whatever, and have whatever working conditions.” [^1] CBS, 60 Minutes
Motivated by these issues, we present a fair pricing framework for the LLM training data market to promote equitable and sustainable generative AI ecosystems. Economic theory suggests that prices should reflect the value delivered to the buyer – signaling quality and alignment [spence1978job]. Guided by this, our framework introduces fairshare pricing based on established data valuation techniques for LLMs [pruthi2020estimating, kwon2023datainf, xia2024less, yu2024mates], which quantify each dataset’s contribution to model performance. Buyers and sellers both have access to data valuation scores, which guide decision-making on both sides: buyers use them to select datasets under budget constraints to maximize utility (i.e., a standard measure of welfare or satisfaction [mas1995microeconomic, von2007theory]), while sellers use them to set prices based on the anticipated demand from buyers. This shared access – enabled by our assumption of information transparency – supports procedural fairness [konovsky2000understanding], improves seller participation, and enhances overall data quality.
@@
Second, we empirically validate our approach through simulations of buyer-seller interactions in data markets. We focus on training open-source LLMs on complex NLP tasks, including math problems [amini2019mathqa], medical diagnosis [jin2020medqa], and physical reasoning [bisk2020piqa]. Analyzing both pricing and valuation outcomes, we find that under fairshare pricing, buyers achieve higher model performance per dollar spent, making it particularly beneficial for those with limited budgets. In addition, our simulations of long-term market dynamics demonstrate that fairshare pricing encourages sustained seller participation, resulting in a stable and sufficient supply of training data over time compared to exploitative pricing. These findings show that our framework’s data-valuation-based pricing not only improves short-term training efficiency, but also ensures the long-term viability of the data market.
Finally, to evaluate the robustness and method-agnostic applicability of our framework, we conduct an ablation study using a diverse set of data valuation methods (including BM25 [trotman2014improvements], $\text{Infl}_{\text{IP}}$ [xia2024less], and Datainf[kwon2023datainf]) – selected for their scalability and efficiency in LLM tasks. Across all variants, our fairshare pricing framework consistently delivers mutually beneficial results for both buyers and sellers in the LLM data market, confirming that its performance is not tied to any specific data valuation technique. This analysis underscores a key strength of our approach: buyers and sellers can flexibly choose valuation methods tailored to their downstream needs without compromising the incentive-aligned structure of the market, and demonstrates the broad applicability of our solution.
@@ 2 Related Work
Data Market Design and Exploitative Pricing. Recent studies on ML data markets focus on platform-based pricing and allocation mechanisms such as auctions and personalized pricing [agarwal2019marketplace, chen2019towards, chen2022selling, yang2022selling, zhang2024survey]. These frameworks typically aim to maximize profit or efficiency, assuming myopic buyers and fixed seller participation. They also consider information transparency ensures that sellers have access to the same data valuation scores used by buyers [ghorbani2020distributional, chen2022selling]. The study was mainly conducted on classic machine learning models rather than large language models.
@@
Concurrently, research has highlighted ethical concerns in LLM data acquisition, particularly around exploitative pricing. Annotation work – central to LLM training – is often outsourced to low-wage workers with minimal labor protections [mason2009financialcrowds, hara2018datadrivenamt, toxtli2021quantifying, remotasks_wapo_2023], who are frequently exposed to harmful content [paul2024bigtech, time2023kenya, weidinger2021ethicalsocialrisksharm]. Several studies document low pay, precarious labor, and lack of protections in global data supply chains [paullada2021datadiscontents, sambasivan2021datawork, gray2019ghost]. Although these studies highlight systemic exploitation in data labor, their findings have yet to be integrated into formal models of LLM data pricing. Economic theory suggests that misaligned compensation undermines efficiency, discourages participation, and lowers data quality. In LLM contexts, this reduces data availability and degrades model performance [techcrunch2024, seetharaman2024gpt5, javed2025aws], underscoring the need for value-aligned pricing mechanisms.
@@
Game-Theoretic Models of Pricing and Participation. In many domains, pricing problems are modeled thorugh game theory. Particularly, those involving decision-making are frequently modeled using Stackelberg games and bilevel optimization [brander1985export]. Although these frameworks have not yet been applied to LLM data market, they have been extensively used in contexts such as resource allocations [bard2013practical], energy market [von2010market, pei2020survey], supply chains [esmaeili2009game], distributed systems [maharjan2013dependable]. In these settings, agents typically optimize revenue, efficiency, or utility while anticipating the strategic responses of others. In dynamic settings, these models extend to repeated interactions, where players balance immediate gains against long-term benefits [stokey1989recursive, brander1985export]. In repeated interactions, fairness perceptions are critical: studies show that perceived pricing unfairness can erode trust and reduce participation [folger1998organizational, folger2001fairness, colquitt2013organizational, adamovic2023organizational].
@@
Data Valuation Methods: Data valuation estimates the contribution of training examples to model performance [hammoudeh2024training]. Examples include influence-function-based methods [koh2017understanding], which estimate data utility via model gradient computations. Variations such of this method enhance efficiency by approximating/bypassing the inverse-Hessian [pruthi2020tracin], or utilizing lower dimension model gradients (e.g., Datainf [kwon2023datainf]), which balance efficiency and accuracy, and make them suitable for the LLM realm. Beyond the influence of loss functions, previous research [wang2022understanding] has examined data valuation with respect to fractional utility functions.
@@
Data valuation approaches are effective for LLM training data selection [xia2024less, yu2024mates, jiao2025datelm] and other applications including toxicity detection [isonuma2024unlearning], memorization analysis [kounavis2023influence], training optimization [van2023hint], and and active sampling [pang2024fairness]. Shapley-based method is another direction [shapley1953value, jia2019towards, ghorbani2019data, ghorbani2020distributional], such as CS-Shaply[schoch2022cs], In-run Shapley[wang2024data], DU-Shapley[garrido2024shapley], etc. Simpler methods like BM25 [trotman2014improvements] offer a model-agnostic baseline based on lexical similarity[akyurek-etal-2022-towards, wu-etal-2024-enhancing-training]. While exact data value estimation is computationally costly at LLM scale, recent studies show that influence-based approximations correlate meaningfully with actual performance outcomes [jiao2024context, wang2024greats]. For example, Jiao et al. [jiao2024context] shows that influence-based methods have high correlation between their estimated data valuation scores and the oracle value, which is obtained through model re-training.
@@ 3.1 Data Market Definition
We formalize the LLM data market by modeling the decision-making processes of sellers and buyers in a non-cooperative Stackelberg game [zhang2024survey, von2010market, maharjan2013dependable]. In this setup, sellers first set prices by anticipating buyer demand. Buyers then respond by selecting datasets to maximize their utility. [^2] Utility is defined as a general economic concept that measures an individual’s welfare, benefit, or satisfaction [mas1995microeconomic, von2007theory], for example, monetary value, model fairness, or other domain-specific benefits. Each player makes decisions based solely on the current state of the market, without considering future updates or interactions.
Notably, our choice of a non-cooperative Stackelberg game fits the LLM data market’s reality: many independent, self-interested transactions where binding grand coalitions are infeasible, such as gig-work platforms [gray2019ghost]. In contrast, cooperative bargaining targets few-party, contractible surplus splits, which is a different regime and question. In this non-cooperative game, we model fairness behaviorally via seller participation/exit that keeps the model tractable, and the leader–follower timing mirrors real price setting and demand response.
@@
Following prior work [tadelis2016reputation], we assume information transparency: all participants observe each dataset’s value. Practically, this can be estimated using data valuation methods that quantify its contribution to LLM performance. These valuation scores – potentially generated by the platform or a trusted third party – serve as common signals informing both pricing and purchasing decisions.
@@
The value $v_{k}$ can be estimated using a data valuation method (e.g., Influence Function [koh2017understanding], DataInf [kwon2023datainf], etc.). Once the data valuation score $v_{k}$ is estimated, the corresponding utility gain $u_{k}$ can be derived as a function of $v_{k}$. We use a task-specific mapping between data value $v_{k}$ and utility $u_{k}:v_{k}\rightarrow\mathbb{R}_{\geq 0}$. Appendix˜C lists common mappings between $u_{k}$ and $v_{k}$ in downstream tasks.
@@ 3.2 Exploitative Pricing: A Lose-Lose Outcome
Assumption˜1 models how the seller respond to exploitative pricing over time. It states that when offered a price below its ideal level, the seller become less likely to stay in the market, consistent with prior studies [gale1994bottom, backus2020sequential]. For any exploitative pricing, the participation probability $P_{T}$ declines multiplicatively, leading to a sustained and irreversible drop in engagement.
@@
Assumption˜2 assumes the buyer values both immediate and future utility gains from acquiring training data. Major LLM developers invest in large-scale data acquisition and model training in expectation of long-term gains in performance, deployment value, and commercial returns [chow2024aws, dotan2024gpt5, li2024time100ai, mayer2025superagency].
@@
Buyer’s Objective. With the previous setup, the buyer aims to maximize expected cumulative utility over an infinite horizon. Let $G(u_{t},b_{t})$ denote this value function, representing the buyer’s expected total cumulative utility at time $t$, conditioned on current utility $u_{t}$ and budget $b_{t}$. Then this objective satisfies the following Bellman equation[bellman1957dynamic]:
@@
In addition to our theoretical findings, we run a simplified dynamic market simulation with one buyer and one seller, making sequential decisions. The utility of the dataset and buyer’s budget varies randomly over time, and seller participation follows $\pi(p_{t},p_{t}^{*})=p_{t}/p_{t}^{*}$. Figures˜2(a) and 2(b) provides consistent evidence that exploitative pricing causes rapid seller exit and an immediate shortage of training data. This dynamic closely resembles the classic “market for lemons” problem, where underpricing drives out high-quality supply, ultimately leading to market collapse [akerlof1978market].
@@ 3.3 Fairshare Pricing: A Win-Win for Sellers and Buyers
And $t^{*}$ is increasing as $\delta$ decreases.
@@ 4.1 Experimental Setup
Buyers and Models. We consider three buyers, each using a standard open-source LLMs: Llama-3.2-Instruct-1b [grattafiori2024llama3herdmodels], Pythia-1b, and Pythia-410m [biderman2023pythia]. These models are pre-trained on different corpora and exhibit varying preferences for downstream post-training data [mai2024fine].
Sellers and Datasets. We focus on challenging, human-annotated tasks: MathQA and GSM8K [amini2019mathqa, cobbe2021training] for math, MedQA [jin2020medqa] for medical diagnosis, and PIQA [bisk2020piqa] for physical reasoning [lu2022survey, liu2023mathematical, ahn2024large]. Table˜1 in Appendix˜F shows dataset splits and examples. We use the training splits as seller data and simulate the market dynamics from Section˜3, treating each task as a market scenario.
@@ 4.2 LLM Data Market Experiments
Exploitative Pricing Leads to Lose-Lose Outcomes: The exploitative pricing method, which sets uniformly low, fixed prices to reflect real-world practices ([cbsnews2024]), offers LLM developers short-term utility gains (see Figures˜4(a) and 4(b)). However, this approach systematically undervalues datasets and unfairly compensates annotators. As a result, sellers exit the market over time, triggering a collapse in data supply. Even well-funded LLM developers are unable to source sufficient data – hindering the advancement of LLMs.
@@ 4.3 Ablation Study: Effect of Different Data Valuation Methods
Data Valuation Methods. We consider the following methods, where each assigns a value to every data sample: (1) Constant baseline – assigns the same value, mimicking flat-rate pricing on platforms; (2) Random baseline – values drawn uniformly from $[0,1]$; (3) Semantic – uses BM25 [trotman2014improvements] to compute average similarity to the representative set; (4) Influence-based: returns a score which leverages learning gradients to estimate a data sample’s avg. contribution to model learning of a representative dataset. Specifically, we use $\text{Infl}_{\text{IP}}$ [pruthi2020tracin, xia2024less] and DataInf  [kwon2023datainf], which are influence-based methods adapted for the LLM realm. See Appendix˜A for additional details.
Furthermore, we adopt a one-step training approach for data valuation, where the value of each data sample is estimated by performing a single training step on it and measuring the resulting change in model performance relative to the original model. Results are presented in in Section˜F.2. This approach serves as an “oracle” for influence-based methods [yu2024mates, pruthi2020tracin], as it directly quantifies data value through model training, although it is more computationally expensive than the previous listed methods.
@@
We also run the error analysis of data valuation methods. Following previous studies [jiao2024context, wang2024greats], we compute the correlation between Oracle and $\text{Infl}_{\text{IP}}$, reporting spearman correlation of $0.54,0.42$ for MathQA and PIQA (see Figure˜8 in Appendix˜F). This underscores the opportunity for future advances in data valuation accuracy and scalability, which can be seamlessly integrated into our flexible fairshare framework.